OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

16 quarters and 30 dimes.

OpenStudy (callisto):

dimes = $0.01? quarter = $0.25?

OpenStudy (anonymous):

what is the work on how to do it

OpenStudy (callisto):

no i'm asking what is dime and what is quarter.. we don't have such currency here :(

OpenStudy (anonymous):

oh quarter is 25cents and a dime is 10

OpenStudy (callisto):

anyway, let x be the no of dimes and y be the no of quarters x+y =46 => x = 46-y -(1) 0.1x+0.25y = 7 -(2) put (1) into (2) 0.1( 46-y)+0.25y = 7 4.6 -0.1y +0.25 = 7 0.15y = 2.4 y=16 therefore x=46-16 =30 therefore there are 30 dimes and 16 quarters

OpenStudy (anonymous):

Thank you so much.. yeah i thought it was 30 dimes but wasn't 100% sure

OpenStudy (callisto):

welcome :)